A box made of steel that is filled with water has a length of 155 meters and a width of 25 meters and on top, a piece of wood is floating that has a mass of 135 million kilograms. What would the depth of the water line be? it also floats in saltwater (density= 1030 kg/m^3)
1
Expert's answer
2020-12-10T11:03:43-0500

"mg=\rho g V\to mg=\rho g l hx\to"


"x=\frac{mg}{\rho g l h}=\frac{135000000\cdot 9.8}{1000\cdot9.8\cdot155\cdot25}\approx35(m)"